MILLIMETER WAVE SPECTROSCOPY OF THE VAN DER WAALS BENDING BANDS OF Ne-HCN

\underline{KENSUKE HARADA}, AKIHITO OKUMURA, TAKEHIKO TANAKA and KEIICHI TANAKA, Department of Chemistry, Faculty of Sciences, Kyushu University, Hakozaki, Higashiku, Fukuoka, 812-8581 JAPAN.

The Ne-HCN cluster is a weakly bound complex with the binding energy (D_{0}) of 39 cm^{−1}.^{a} In the present work, we have observed the vdW bending bands of Ne-HCN. A pulsed jet millimeter-wave spectrometer was used for the experiment. The millimeter wave radiation passed twenty times through the jet expansion using a White type optical path to attain high sensitivity. The 14 and 22 vdW transitions were observed for the Σ_{1}-Σ_{0} and Π_{1}-Σ_{0} bands in the frequency region of 79-201 GHz. Most lines were split into the hyperfine components of nitrogen nucleus. From the eQq constants, the root mean square amplitudes ⟨θ^{2}⟩^{1/2} of the HCN part were estimated to be 47^{∘}, 56^{∘}, and 63^{∘} for the Σ_{0}, Σ_{1}, and Π_{1} states. The center of mass distance of the intermolecular bond were derived from the rotational constants to be 3.89, 3.85, and 3.66 Å for those states.

aG. Murdachaew, A. J. Misquitta, R. Bukowski, and K. Szalewicz, J. Chem. Phys. 114, 764(2001).